jhh:

Using the graph of f(x) = log10x below, approximate the value of y in the equation 10y = 5.

Vocaloid:

Presuming it's 10^y = 5 not 10y = 5 Anyway, our graph is y = log_10(x) raising both sides to base 10 gives us 10^y = x now, comparing this to 10^y = 5, this simply means you need to find the value of y when x = 5